# arxiv-appendix

### v1

@@ 2.3.5 Refining Bins
The algorithm offers several advantages. First, it is tuning-free, meaning that no hyperparameter needs to be predetermined. It also does not require knowing the total target sample size $n_{\mathrm{P}}$, but only a rough estimate of its order, which is used for calculating $\log(n_{\mathrm{P}})$. Recall that $\log(n_{\mathrm{P}})$ is used for the determination of $t_{s}^{j}\geq\log^{2}(n_{\mathrm{P}})$ as well as the constant $C_{n_{\mathrm{P}}}$ in (14). This makes the algorithm suitable for online deployment where the future sample size can be estimated only approximately. Additionally, it is sequentially-interactive: once a user sends their message, $\tilde{V}_{i,k,s}^{\mathrm{P},j}$ and $\tilde{U}_{i,k,s}^{\mathrm{P},j}$, it can safely exit the system. This property is particularly beneficial in industrial settings, since it is challenging to keep users engaged continuously. Finally, as shown previously in Section 2.2, the algorithm also has theoretical optimality.
We highlight the key differences between our policy and those proposed in Perchet and Rigollet (2013) and Cai et al. (2024). Perchet and Rigollet (2013) also employ successive elimination-type algorithms, while Cai et al. (2024) utilize an algorithm with carefully designed exploration and exploitation phases. Both approaches assume that each bin has access to all historical information, which is not feasible under the LDP constraints. For instance, suppose that we query the $i$-th sample for $\tilde{U}_{i,k,s}^{\mathrm{P},j}$ with privacy budget $\varepsilon$. If $B_{s}^{j}$ is subsequently split, this sample cannot be used for either $\tilde{U}_{i,k,s+1}^{\mathrm{P},2j-1}$ or $\tilde{U}_{i,k,s+1}^{\mathrm{P},2j}$ due to LDP constraints. More specifically, due to privatization, we lose the ability to determine which sub-bin the sample belongs to, rendering the $i$-th user’s data unusable in the subsequent learning process. This is also the reason for designing the indicators $a_{i,s}^{j}$, which is to disable the past data after a bin is refined. One might suggest querying the sample multiple times using privacy composition techniques (e.g. Dwork et al., 2010; Kairouz et al., 2015). However, this approach would require dividing the already limited privacy budget $\varepsilon$, yielding a loss of efficiency. Another alternative would be to create a fixed partition with a pre-determined depth. Though the fixed partition collects information from all samples, it introduces a highly sensitive hyperparameter, i.e. the depth. Our mechanism is efficient in terms of the multiplicative constant before the added noise and does not require additional hyperparameters.
@@ 3 Auxiliary Data Source: A Jump-start
In this section, we further consider transfer learning and discuss how auxiliary datasets can bring a jump-start effect to the bandit problem.
@@ 3.1 Preliminaries
In addition to $\{\tilde{Z}^{\mathrm{P}}_{t}\}_{t\geq 1}$ coming in sequentially, we assume that there are $M\in\mathbb{Z}_{+}$ auxiliary data sources $\mathcal{D}^{\mathrm{Q}_{1}},\dots,\mathcal{D}^{\mathrm{Q}_{M}}$, where $\mathcal{D}^{\mathrm{Q}_{m}}\coloneqq\{Z^{\mathrm{Q}_{m}}_{i}\}_{i=1}^{n_{\mathrm{Q}_{m}}}$ and $Z^{\mathrm{Q}_{m}}_{i}=(X_{i}^{\mathrm{Q}_{m}},\pi_{i}^{\mathrm{Q}_{m}}(X_{i}^{\mathrm{Q}_{m}})$, $Y_{i}^{\mathrm{Q}_{m},(\pi_{i}^{\mathrm{Q}_{m}}(X_{i}^{\mathrm{Q}_{m}}))})$ are generated similarly on $\mathcal{X}\times[0,1]^{K}$ using policy $\pi^{\mathrm{Q}_{m}}$. For now, we assume that these data sources are historical datasets, meaning that all auxiliary data are ready to be queried before initiating interaction with the target data. We discuss in Section 5 the case where the auxiliary data are in the form of streaming data - a scenario conforming to the multi-task learning setting. We assume $\pi_{i}^{\mathrm{Q}_{m}}$’s are fixed behavior policies. Behavior policy is suitable for describing batched data (Lange et al., 2012; Levine et al., 2020) and is used in the literature of MAB with auxiliary data (e.g. Zhang and Bareinboim, 2017; Cai et al., 2024).
Distribution shift.
We allow differences between the distributions of target and auxiliary data by adopting the covariate shift setting. In particular, we allow the marginal distributions of covariates in the $\mathrm{P}$-bandit and $\mathrm{Q}$-bandits to be different (i.e. $\mathrm{P}_{X}\neq\mathrm{Q}_{m,X}$, for all $1\leq m\leq M$), while the distributions of rewards conditioned on the covariate are assumed to be identical, i.e. $\mathrm{P}_{Y^{(k)}|X}=\mathrm{Q}_{m,Y^{(k)}|X}$ for all $1\leq k\leq K$ and $1\leq m\leq M$. We denote the common reward function of the $k$-th arm as $f_{k}(x)\coloneqq f_{k}^{\mathrm{P}}(x)\equiv f_{k}^{\mathrm{Q}_{m}}(x)$ for all $k\in[K]$ and $x\in\mathcal{X}$.
Privacy.
We consider the case where policy $\pi$ receives information from $\mathcal{D}^{\mathrm{Q}_{m}}$ via a sequentially-interactive $\varepsilon_{m}$-LDP mechanism. The privacy budgets $\varepsilon_{m}$ are allowed to vary across the $M$ datasets. We define the class of policies that are $(\varepsilon,\varepsilon_{1},\dots,\varepsilon_{M})$-LDP with respect to $(\mathcal{D}^{\mathrm{P}},\mathcal{D}^{\mathrm{Q}_{1}},\dots,\mathcal{D}^{\mathrm{Q}_{M}})$ by $\Pi(\varepsilon,\varepsilon_{1},\ldots,\varepsilon_{M})$.
3.2 Minimax Regret Bound
We characterize the connections and differences between the auxiliary and target distributions through the following assumptions.
@@ 3.2 Minimax Regret Bound
$\mathrm{Q}_{mX}(B(x,r))\geq C_{\gamma_{m}}r^{\gamma_{m}}\mathrm{P}_{X}(B(x,r)),\quad\forall x\in\mathcal{X},r\in(0,1],$ (18)
@@
Given Definition 3.1 and 3.2, we consider the following class of contextual MABs
@@
We comment on these concepts. The transfer exponent is a widely used concept for quantifying covariate shift (e.g. Kpotufe and Martinet, 2021; Cai et al., 2024). It requires that the minimum probability under $\mathrm{Q}$ within a given ball is comparable to that under $\mathrm{P}$. Clearly, if $\mathrm{Q}_{m}=\mathrm{P}$, then $\gamma_{m}=0$. A larger $\gamma_{m}$ indicates a greater distribution discrepancy. Definition 3.2 pertains to the historical data setting, suggesting that the behavior policies should sufficiently explore all arms.
@@
$+$ $\sum_{m=1}^{M}\left\{\frac{\kappa_{m}^{2}n_{\mathrm{Q}_{m}}}{K^{2}}(e^{\varepsilon_{m}}-1)^{2}\right\}^{\frac{2+2d}{2+2d+2\gamma_{m}}}\wedge\left(\frac{\kappa_{m}n_{\mathrm{Q}_{m}}}{K}\right)^{\frac{2+2d}{2+d+\gamma_{m}}}\Bigg{]}^{-\frac{1+\beta}{2+2d}},$ (21)
where $c>0$ is a constant depending only on $d,C_{L},\beta,M,\gamma$.
In the lower bound (21), the terms associated with the auxiliary data contain a factor of $K$, while the term associated with the target data not. This arises from our assumption that the source policies are behavior policies, which simplifies the theoretical analysis. When compared to the upcoming upper bound (22), we observe that these dependencies on $K$ are tight. Specifically, for the non-private and private terms, the dependencies are $K$ and $K^{2}$, respectively. This indicates that increasing the number of arms introduces greater challenges under privacy constraints, echoing our discussions at the end of Section 2.2.
@@
Consider the class of distributions $\Lambda(K,\beta,\gamma,\kappa)$ defined in (20) and the class of LDP policies $\Pi(\varepsilon,\varepsilon_{1},\ldots,\varepsilon_{M})$. Suppose that $(\mathrm{P},\{\mathrm{Q}_{m}\}_{m=1}^{M})\in\Lambda(K,\beta,\gamma,\kappa)$. Then, we have $(1)$ the policy $\pi$ given by Algorithm 3 satisfies $\pi\in\Pi(\varepsilon,\varepsilon_{1},\ldots,\varepsilon_{M})$ and $(2)$ with probability at least $1-n^{-2}$, the regret of $\pi$ has
$R_{n_{\mathrm{P}}}(\pi)\leq Cn_{\mathrm{P}}\Bigg{[}$$\left\{\frac{n_{\mathrm{P}}\varepsilon^{2}}{K^{2}\log(n)}\right\}\wedge\left(\frac{n_{\mathrm{P}}}{K\log(n)}\right)^{\frac{2+2d}{2+d}}$
$+$ $\sum_{m=1}^{M}\left\{\frac{\kappa_{m}^{2}n_{\mathrm{Q}_{m}}\varepsilon_{m}^{2}}{K^{2}\log(n)}\right\}^{\frac{2+2d}{2+2d+2\gamma_{m}}}\wedge\left\{\frac{\kappa_{m}n_{\mathrm{Q}_{m}}}{K\log(n)}\right\}^{\frac{2+2d}{2+d+\gamma_{m}}}\Bigg{]}^{-\frac{1+\beta}{2+2d}},$ (22)
where $C>0$ is a constant depending only on $d,C_{L},\beta,M,\gamma$.
Treating the number of arms $K$ as an absolute constant, considering a challenging, high-privacy regime that $\varepsilon,\max_{m=1}^{M}\varepsilon_{m}\in(0,1]$, we have that, up to the logarithmic factors, the minimax rate of the regret is of order
@@
Compared to (7), we observe that (23) has an increased effective sample size, showing the superiority of including auxiliary data. The contributions of the auxiliary data, compared to target data, are reduced by a polynomial factor of $\kappa_{m}$ and an exponential factor of $\gamma_{m}$, for the auxiliary dataset indexed by $m\in[M]$. When $\kappa_{m}$ is small, there are arms rarely explored, which could potentially be the best arm, thereby limiting the contributions of the auxiliary datasets. When $\gamma_{m}$ is large, the marginal distribution $\mathrm{Q_{m,X}}$ can deviate significantly from $\mathrm{P}_{X}$, providing redundant information in regions where it is unnecessary. This also reduces the effective sample size of $\mathcal{D}^{\mathrm{Q}_{m}}$.
3.3 Methodology
We now demonstrate how to leverage the auxiliary dataset to enhance the performance of our estimator by designing an additional jump-start stage, where we apply a similar successive elimination procedure starting from $n_{\mathrm{Q}_{1}}$ samples of $\mathcal{D}^{\mathrm{Q}_{1}}$, continuing with $n_{\mathrm{Q}_{2}}$ samples of $\mathcal{D}^{\mathrm{Q}_{2}}$, finishing off with the $n_{\mathrm{Q}_{M}}$ samples of $\mathcal{D}^{\mathrm{Q}_{M}}$. Each sample interacts with the policy only once. We then proceed with learning on $\mathcal{D}^{\mathrm{P}}$. Therefore, learning on the target data can utilize the refined partition and the set of the selected active arms learned via the source datasets. For an illustration of the benefits of this manner, see Figure 6 in the experiments section.
We simplify the notation by re-indexing the time indices in each dataset with $t\in[n_{\mathrm{P}}+\sum_{m=1}^{M}n_{\mathrm{Q}_{m}}]$, defined as the total number of samples that have interacted with policy. Since the datasets interact sequentially, we define
@@ 3.3 Methodology
for all $m\in[M]$, representing the time index before, during, and after the interaction of $m$-th dataset. Analogously, define the target time index by $T_{0}(t)=(t-\sum_{m=1}^{M}n_{\mathrm{Q}_{m}})\vee 0$. For notational simplicity, denote $\mathrm{P}$ as $\mathrm{Q}_{0}$ and write $n_{\mathrm{Q}_{0}}=n_{\mathrm{P}}$, $\varepsilon_{0}=\varepsilon$. For $m\in[M]\cup\{0\}$, define $a_{i,s}^{m,j}=1$ if $Z_{i}^{\mathrm{Q}_{m}}$ is used to update $B_{s}^{j}$ and 00 otherwise. Let the cumulatively used sample size be $t_{s}^{m,j}=\sum_{i=1}^{T_{m}(t)}a_{i,s}^{m,j}$ and the maximum sample size $n=n_{\mathrm{P}}\vee(\max_{m=1}^{M}n_{\mathrm{Q}_{m}})$. Similar to (11) and (12), we encode the information from the auxiliary data by
${V}_{T_{m}(t),k,s}^{\mathrm{Q}_{m},j}=$ $Y_{T_{m}(t)}^{\mathrm{Q}_{m},(\pi_{T_{m}(t)}^{\mathrm{Q}_{m}}(X_{T_{m}(t)}^{\mathrm{Q}_{m}}))}\boldsymbol{1}(X_{T_{m}(t)}^{\mathrm{Q}_{m}}\in B^{j}_{s})\boldsymbol{1}(\pi_{T_{m}(t)}^{\mathrm{Q}_{m}}(X_{T_{m}(t)}^{\mathrm{Q}_{m}})=k),$ (24)
${U}_{T_{m}(t),k,s}^{\mathrm{Q}_{m},j}=$ $\boldsymbol{1}(X_{T_{m}(t)}^{\mathrm{Q}_{m}}\in B^{j}_{s})\boldsymbol{1}(\pi_{T_{m}(t)}^{\mathrm{Q}_{m}}(X_{T_{m}(t)}^{\mathrm{Q}_{m}})=k),$
for $t\in[\sum_{m=1}^{M}n_{\mathrm{Q}_{m}}],k\in[K],B_{s}^{j}\in\mathcal{B}_{t},m\in[M]$. They are then privatized as
@@
We present the detailed algorithm for leveraging auxiliary data in Algorithm 3. The algorithm essentially repeats the sequential procedures outlined in Algorithm 2 on the auxiliary data before interacting with the target data. Unlike the target data, the auxiliary datasets already contain executed policies $\pi_{T_{m}(t)}^{\mathrm{Q}_{m}}(X_{T_{m}(t)}^{\mathrm{Q}_{m}})$. As a result, learning on the auxiliary data does not involve making instant decisions based on the learned policy. However, encoding (24) still requires the active partition $\mathcal{B}_{t}$, which is gradually updated throughout the algorithm. Since multiple datasets are involved, Algorithm 3 also requires a multiple-source version of the local estimator and confidence bound. These are defined as follows.
Considering the local estimates, it is likely that several datasets will contribute to the same bin. Their contributions need to be balanced due to different variance levels induced by the LDP constraints. Without LDP constraints, such weighting is unnecessary (e.g. Cai et al., 2024) and we will elaborate the connection later. We define the multiple-source estimator as
@@
In (26), the influence of each data is controlled by the weight $\lambda_{t,k,s}^{m,j}$. Specifically, we let
@@
$\lambda_{t,k,s}^{m,j}=\left|\frac{\varepsilon_{m}^{2}}{t_{s}^{m,j}}\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}\right|\wedge\boldsymbol{1}\left\{t_{s}^{m,j}\geq\log^{2}(n)\right\}.$ (27)
@@
Here, the condition $\boldsymbol{1}\{t_{s}^{m,j}\geq\log^{2}(n)\}$ ensures that the $m$-th dataset has provided sufficient samples, a requirement by the theoretical validity of our confidence bound (28). Without the condition, the weight is zero, and the $m$-th data is excluded from $\tilde{f}^{t}_{k}(x)$. When the condition is met, the weight $\lambda_{t,k,s}^{m,j}$ depends on two factors that characterize the information from the $m$-th dataset. One is $({t_{s}^{m,j}})^{-1}\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}$, which approximates the number of samples within the bin that has pulled arm $k$ and represents the quantity of information. The other is the privacy budget $\varepsilon_{m}$, which reflects the accuracy of each $\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}$ and represents the quality of information. If both factors are relatively large, the dataset is considered informative and is therefore assigned a large weight. In the non-private case (i.e. $\varepsilon_{m}=\infty$), our choice of $\lambda$ indicates that all weights are assigned equal values of $1$, which is consistent with Cai et al. (2024).
Following the same reasoning as in Section 2.3.4, we define

### v2

@@ 2.3.6 Summary and discussions
Our upper bound algorithm offers several advantages. First, it is essentially tuning-free, meaning that no hyperparameter needs to be predetermined. Moreover, it is sequentially-interactive: once a user sends the privatized $\tilde{V}_{i,k,s}^{\mathrm{P},j}$ and $\tilde{U}_{i,k,s}^{\mathrm{P},j}$, it can safely exit the system (e.g. websites). This property is particularly beneficial in industrial settings since it is challenging to continuously track and communicate with users once they leave the system. Finally, as shown previously in Section 2.2, our algorithm achieves the near-optimal regret upper bound.
As discussed before, our algorithm is inspired by the adaptive binning and successive elimination (ABSE) algorithm proposed in Perchet and Rigollet, (2013). Here, we highlight their key differences, which stems from the LDP constraints. First, to preserve privacy, our algorithm separates the user-server operations and only allows privatized information exchange between the two sides. Therefore, it is necessary to design new and efficient private nonparametric reward function estimator and the corresponding confidence bound for our policy, which is more challenging than the non-private setting. Second, without privacy concerns, ABSE has the luxury of being able to access and thus leverage all past information for updating its policy, which is not feasible under the LDP constraints. As a concrete example, suppose that at time $i$, for a given bin $B_{s}^{j}$ in the active partition $\mathcal{B}_{i}$, we query the $i$-th user with a privacy budget $\varepsilon$ to construct $\tilde{U}_{i,k,s}^{\mathrm{P},j}$. If $B_{s}^{j}$ is subsequently refined into sub-bins $B_{s+1}^{2j-1}$ and $B_{s+1}^{2j}$, the raw data of the $i$-th user cannot be re-queried to construct $\tilde{U}_{i,k,s+1}^{\mathrm{P},2j-1}$ or $\tilde{U}_{i,k,s+1}^{\mathrm{P},2j}$ as we have used up the $\varepsilon$ privacy budget. In addition, due to privatization, the $\tilde{U}_{i,k,s}^{\mathrm{P},j}$ quantity cannot be utilized via post-processing to (approximately) determine which sub-bin the $i$-th user belongs to, rendering it unusable in the subsequent learning process. Indeed, this is why our algorithm designs the indicators $a_{i,s}^{j}$, which disables past (privatized) information once a bin is refined.
One might suggest querying a user multiple times using privacy composition techniques (e.g. Dwork et al.,, 2010). However, this approach would require dividing the already limited privacy budget $\varepsilon$, yielding a loss of efficiency. Moreover, it requires to continuously track and communicate with the users, which is not ideal under industry settings. Another option would be to create a fixed partition with a pre-determined depth. Though the fixed partition can collect (privatized) information from all samples, it introduces a highly sensitive hyperparameter, i.e. the depth of the partition, the choice of which is not obvious and thus is undesirable in practice.
@@ 3 Auxiliary Data Source: A Jump-start
In this section, we further extend our study to transfer learning (TL) and discuss how auxiliary data can bring a jump-start effect to the nonparametric contextual MAB under the LDP constraints.
@@ 3.1 Preliminaries
In addition to the target data $\{\tilde{Z}^{\mathrm{P}}_{t}\}_{t\geq 1}$, which comes in sequentially, we assume that there are $M\in\mathbb{Z}_{+}$ auxiliary datasets $\mathcal{D}^{\mathrm{Q}_{1}},\dots,\mathcal{D}^{\mathrm{Q}_{M}}$, where $\mathcal{D}^{\mathrm{Q}_{m}}\coloneqq\{Z^{\mathrm{Q}_{m}}_{i}\}_{i=1}^{n_{\mathrm{Q}_{m}}}$ and $Z^{\mathrm{Q}_{m}}_{i}=(X_{i}^{\mathrm{Q}_{m}},\pi_{i}^{\mathrm{Q}_{m}}(X_{i}^{\mathrm{Q}_{m}})$, $Y_{i}^{\mathrm{Q}_{m},(\pi_{i}^{\mathrm{Q}_{m}}(X_{i}^{\mathrm{Q}_{m}}))})$ are generated similarly on $\mathcal{X}\times[0,1]^{K}$ based on policy $\pi^{\mathrm{Q}_{m}}$. For now, we assume that the auxiliary data are historical datasets, meaning that all auxiliary data are ready to be queried before we initiate interaction with the target data. We discuss in Section 5 the case where the auxiliary data are in the form of streaming data - a scenario conforming to the multi-task learning setting. We assume $\pi_{i}^{\mathrm{Q}_{m}}$’s are fixed behavior policies, i.e. $\pi^{\mathrm{Q}_{m}}_{i}\equiv\pi^{\mathrm{Q}_{m}}$, $m\in[M]$ and $i\in[n_{\mathrm{Q}_{m}}]$. Behavior policy is suitable for describing batched data (Lange et al.,, 2012; Levine et al.,, 2020) and is widely used in the literature of MAB with auxiliary data (e.g. Zhang and Bareinboim,, 2017; Cai et al.,, 2024).
Distribution shift. We allow differences between the distributions of target and auxiliary data by adopting the covariate shift setting. In particular, we allow the marginal distributions of covariates in the $\mathrm{P}$-bandit and $\mathrm{Q}$-bandits to be different (i.e. $\mathrm{P}_{X}\neq\mathrm{Q}_{m,X}$, for all $1\leq m\leq M$), while the distributions of rewards conditioned on the covariate are assumed to be identical, i.e. $\mathrm{P}_{Y^{(k)}|X}=\mathrm{Q}_{m,Y^{(k)}|X}$ for all $1\leq k\leq K$ and $1\leq m\leq M$. We denote the common reward function of the $k$-th arm as $f_{k}(x)\coloneqq f_{k}^{\mathrm{P}}(x)\equiv f_{k}^{\mathrm{Q}_{m}}(x)$ for all $k\in[K]$ and $x\in\mathcal{X}$.
Privacy. We allow the target data policy $\pi$ to receive information from $\mathcal{D}^{\mathrm{Q}_{m}}$ via a sequentially-interactive $\varepsilon_{m}$-LDP mechanism. The privacy budgets $\varepsilon_{m}$ are allowed to vary across the $M$ auxiliary datasets. We denote the class of policies that are $(\varepsilon,\varepsilon_{1},\dots,\varepsilon_{M})$-LDP with respect to $(\mathcal{D}^{\mathrm{P}},\mathcal{D}^{\mathrm{Q}_{1}},\dots,\mathcal{D}^{\mathrm{Q}_{M}})$ by $\Pi(\varepsilon,\varepsilon_{1},\ldots,\varepsilon_{M})$.
3.2 Minimax Optimal Regret Bound
We first characterize the connections and differences between the auxiliary and target distributions through the following assumptions.
@@ 3.2 Minimax Optimal Regret Bound
$\mathrm{Q}_{m,X}(B(x,r))\geq C_{\gamma_{m}}r^{\gamma_{m}}\mathrm{P}_{X}(B(x,r)),\quad\forall x\in\mathcal{X},r\in(0,1],$ (18)
@@
Given Definitions 3.1 and 3.2, we consider the following class of contextual MABs
@@
We comment on these concepts. The transfer exponent is a widely used term for quantifying covariate shift (e.g. Kpotufe and Martinet,, 2021; Cai et al.,, 2024). It requires that the minimum probability under $\mathrm{Q}$ within a given ball is comparable to that under $\mathrm{P}$. Clearly, if $\mathrm{Q}_{m}=\mathrm{P}$, then $\gamma_{m}=0$. A larger $\gamma_{m}$ indicates a greater distribution discrepancy. Definition 3.2 pertains to the historical data setting, suggesting that the behavior policies should sufficiently explore all arms.
@@
$+$ $\sum_{m=1}^{M}\left(\frac{\kappa_{m}^{2}n_{\mathrm{Q}_{m}}}{K^{2}}(e^{\varepsilon_{m}}-1)^{2}\right)^{\frac{2+2d}{2+2d+2\gamma_{m}}}\wedge\left(\frac{\kappa_{m}n_{\mathrm{Q}_{m}}}{K}\right)^{\frac{2+2d}{2+d+\gamma_{m}}}\Bigg{]}^{-\frac{1+\beta}{2+2d}},$ (21)
where $c>0$ is an absolute constant depending only on $d,C_{L},\beta,M,\gamma$.
Theorem 3.3 indicates that the regret can be improved when auxiliary data is available and it further recovers the lower bound result in Theorem 2.5 when setting $M=0$. In the lower bound (21), the term associated with the auxiliary data contains a factor of $K$, while the term associated with the target data does not. This arises from our assumption that the policies that generate the auxiliary data are fixed behavior policies (i.e. not adaptively updated over time). In addition, note that for the term associated with the auxiliary data in (21), the dependencies on the number of arms are $K$ and $K^{2}$ for its non-private and private components, respectively, suggesting that increasing the number of arms introduces greater challenges under privacy constraints.
@@
Consider the class of distributions $\Lambda(K,\beta,\gamma,\kappa)$ defined in (20) and the class of LDP policies $\Pi(\varepsilon,\varepsilon_{1},\ldots,\varepsilon_{M})$. Suppose that $(\mathrm{P},\{\mathrm{Q}_{m}\}_{m=1}^{M})\in\Lambda(K,\beta,\gamma,\kappa)$. Then, we have that the policy $\pi$ given by Algorithm 3 satisfies $\pi\in\Pi(\varepsilon,\varepsilon_{1},\ldots,\varepsilon_{M})$ and with probability at least $1-n^{-2}$, the regret of $\pi$ satisfies that
$R_{n_{\mathrm{P}}}(\pi)\leq Cn_{\mathrm{P}}\Bigg{[}$ $\left(\frac{n_{\mathrm{P}}\varepsilon^{2}}{K^{2}\log(n)}\right)\wedge\left(\frac{n_{\mathrm{P}}}{K\log(n)}\right)^{\frac{2+2d}{2+d}}$
$+$ $\sum_{m=1}^{M}\left(\frac{\kappa_{m}^{2}n_{\mathrm{Q}_{m}}\varepsilon_{m}^{2}}{K^{2}\log(n)}\right)^{\frac{2+2d}{2+2d+2\gamma_{m}}}\wedge\left(\frac{\kappa_{m}n_{\mathrm{Q}_{m}}}{K\log(n)}\right)^{\frac{2+2d}{2+d+\gamma_{m}}}\Bigg{]}^{-\frac{1+\beta}{2+2d}},$ (22)
where $C>0$ is an absolute constant depending only on $d,C_{L},\beta,M,\gamma$ and $n=n_{\mathrm{P}}\vee(\max_{m=1}^{M}n_{\mathrm{Q}_{m}})$ is the maximum sample size.
Treating the number of arms $K$ as a constant and considering the challenging, high-privacy regime that $\max\{\varepsilon,\varepsilon_{1},\cdots,\varepsilon_{M}\}\in(0,1]$, we have that, up to the logarithmic factors, the minimax rate of the regret is of order
@@
Compared to the minimax rate without TL in (7), we observe that (23) has an increased effective sample size, showing the benefit of auxiliary data. The contributions of the auxiliary data, compared to target data, are reduced by a polynomial factor of $\kappa_{m}$ and an exponential factor of $\gamma_{m}$, which is indeed intuitive and interpretable. When $\kappa_{m}$ is small, there are arms rarely explored, which could potentially be the best arm, thereby limiting the contributions of the auxiliary datasets. When $\gamma_{m}$ is large, the marginal distribution $\mathrm{Q}_{m,X}$ can deviate significantly from $\mathrm{P}_{X}$, providing redundant information in regions where it is unnecessary. This also reduces the effective sample size of $\mathcal{D}^{\mathrm{Q}_{m}}$.
3.3 Upper Bound Methodology
We now demonstrate how to leverage the auxiliary data to enhance the performance of our policy in Algorithm 2 by designing an additional jump-start stage, where we apply a similar arm elimination procedure starting from $n_{\mathrm{Q}_{1}}$ samples of $\mathcal{D}^{\mathrm{Q}_{1}}$, continuing with $n_{\mathrm{Q}_{2}}$ samples of $\mathcal{D}^{\mathrm{Q}_{2}}$, and finishing off with the $n_{\mathrm{Q}_{M}}$ samples of $\mathcal{D}^{\mathrm{Q}_{M}}$. Each sample interacts with the policy only once. We then proceed with learning on $\mathcal{D}^{\mathrm{P}}$. Therefore, learning on the target data can utilize the refined partition and the set of the selected active arms learned via the source data. For a concrete illustration of such benefits, see Figure 6 in the numerical experiments section.
We proceed by defining some necessary notations. First, we simplify the notation by re-indexing the time indices in each dataset with $t\in[n_{\mathrm{P}}+\sum_{m=1}^{M}n_{\mathrm{Q}_{m}}]$, defined as the total number of users that have interacted with policy $\pi$. We further define
@@ 3.3 Upper Bound Methodology
for all $m\in[M]$, which gives the total number of users from the $m$-th auxiliary dataset that have interacted with policy $\pi$ up to time $t.$ Analogously, define the target time index by $T_{0}(t)=(t-\sum_{m=1}^{M}n_{\mathrm{Q}_{m}})\vee 0$. For notational simplicity, we further denote $\mathrm{P}$ as $\mathrm{Q}_{0}$ and write $n_{\mathrm{Q}_{0}}=n_{\mathrm{P}}$, $\varepsilon_{0}=\varepsilon$. For $m\in[M]\cup\{0\}$, define $a_{i,s}^{m,j}=1$ if $Z_{i}^{\mathrm{Q}_{m}}$ is used to update $B_{s}^{j}$ and 00 otherwise. Let the cumulative sample size be $t_{s}^{m,j}=\sum_{i=1}^{T_{m}(t)}a_{i,s}^{m,j}$. Similar to (11) and (12), we encode the information from the auxiliary data by
${V}_{T_{m}(t),k,s}^{\mathrm{Q}_{m},j}=$ $Y_{T_{m}(t)}^{\mathrm{Q}_{m},\big{(}\pi_{T_{m}(t)}^{\mathrm{Q}_{m}}(X_{T_{m}(t)}^{\mathrm{Q}_{m}})\big{)}}\boldsymbol{1}(X_{T_{m}(t)}^{\mathrm{Q}_{m}}\in B^{j}_{s})\boldsymbol{1}(\pi^{\mathrm{Q}_{m}}(X_{T_{m}(t)}^{\mathrm{Q}_{m}})=k),$ (24)
${U}_{T_{m}(t),k,s}^{\mathrm{Q}_{m},j}=$ $\boldsymbol{1}(X_{T_{m}(t)}^{\mathrm{Q}_{m}}\in B^{j}_{s})\boldsymbol{1}(\pi^{\mathrm{Q}_{m}}(X_{T_{m}(t)}^{\mathrm{Q}_{m}})=k),$
for $t\in[\sum_{m=1}^{M}n_{\mathrm{Q}_{m}}],k\in[K],B_{s}^{j}\in\mathcal{B}_{t}$ and $m\in[M]$. They are then privatized as
@@
We present the detailed algorithm for leveraging auxiliary data in Algorithm 3. The algorithm essentially repeats the sequential procedures outlined in Algorithm 2 on the auxiliary data before interacting with the target data. Unlike the target data, the auxiliary datasets already contain executed policies $\pi^{\mathrm{Q}_{m}}(X_{T_{m}(t)}^{\mathrm{Q}_{m}})$. As a result, learning on the auxiliary data does not involve making instant decisions based on the learned policy $\pi$. However, the active partition $\mathcal{B}_{t}$ and the associated active arms sets are gradually updated throughout the interaction with auxiliary data.
Importantly, since multiple datasets are involved, Algorithm 3 requires a multiple-source version of the local estimator and confidence bound for the reward function. In particular, it is likely that several datasets may contribute to the local estimates of the same bin. Thus, to achieve optimal estimation efficiency, their contributions need to be carefully weighted due to different variance levels induced by the LDP constraints. To this end, we propose a novel multiple-source local estimator where
@@
In (26), the influence of each dataset is controlled by the weight $\lambda_{t,k,s}^{m,j}$. Specifically, we set
@@
$\lambda_{t,k,s}^{m,j}=\left|\frac{\varepsilon_{m}^{2}}{t_{s}^{m,j}}\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}\right|\wedge\boldsymbol{1}\left\{t_{s}^{m,j}\geq\log^{2}(n)\right\},$ (27)
@@
where recall we denote $n=n_{\mathrm{P}}\vee(\max_{m=1}^{M}n_{\mathrm{Q}_{m}}).$ Here, the condition $\boldsymbol{1}\{t_{s}^{m,j}\geq\log^{2}(n)\}$ ensures that the $m$-th dataset has provided sufficient samples, a requirement needed for the theoretical validity of our confidence bound in (28). When the condition is unmet, the weight is zero, and the $m$-th data is excluded from $\tilde{f}^{t}_{k}(x)$. When the condition holds, the weight $\lambda_{t,k,s}^{m,j}$ depends on two factors that characterize the information from the $m$-th dataset. One is $({t_{s}^{m,j}})^{-1}\sum_{i=1}^{T_{m}(t)}\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}a_{i,s}^{m,j}$, which approximates the proportion of samples within the bin that pulled arm $k$ and represents the quantity of information. The other is the privacy budget $\varepsilon_{m}$, which reflects the accuracy of each $\tilde{U}_{i,k,s}^{\mathrm{Q}_{m},j}$ and represents the quality of information. If both factors are relatively large, the dataset is considered informative and is therefore assigned a large weight. We note that without LDP constraints, such weighting scheme is not necessary. Indeed, in the non-private case (i.e. $\varepsilon_{m}=\infty$), our choice of $\lambda$ indicates that all weights are assigned equal values of $1$, which is consistent with non-private transfer learning for nonparametric contextual MAB (Cai et al.,, 2024).
Moreover, we define the corresponding confidence bound as